# Fractal Graph Contrastive Learning

### v2

@@ Appendix A Proofs
For each candidate $H_{i}$, the single-sample fractal loss is
$\ell_{\mathrm{fractal}}(G,H_{i})=-\frac{s_{i}}{\tau}-\log w_{i}+\log Z,\qquad i=1,2,$
where $Z$ is the common partition term. Taking the difference,
$\ell_{\mathrm{fractal}}(G,H_{2})-\ell_{\mathrm{fractal}}(G,H_{1})=-\frac{s_{2}-s_{1}}{\tau}-\log\frac{w_{2}}{w_{1}}.$
Since $\Delta_{2}>\Delta_{1}$ and $w_{i}=\exp(\alpha\Delta_{i})$, we have
$\log\frac{w_{2}}{w_{1}}=\alpha(\Delta_{2}-\Delta_{1})>0.$
By the assumption $s_{1}-s_{2}\leq\varepsilon=\tau\alpha(\Delta_{2}-\Delta_{1})$, it follows that
$-\frac{s_{2}-s_{1}}{\tau}\leq\alpha(\Delta_{2}-\Delta_{1}),$
hence the difference of losses is negative:
$\ell_{\mathrm{fractal}}(G,H_{2})-\ell_{\mathrm{fractal}}(G,H_{1})<0.$
Therefore $\ell_{\mathrm{fractal}}(G,H_{2})<\ell_{\mathrm{fractal}}(G,H_{1})$, completing the proof. ∎
@@
Lower bound $\boldsymbol{\Omega(V^{2})}$. Consider a path of $V$ vertices. At the smallest scale (covering radius $1$) a single box covers at most two vertices, so roughly $V/2$ boxes must be chosen. Each choice is made by scanning all currently uncovered vertices to find the one whose radius-$1$ neighbourhood is largest: first $V$ scans, then $V-1$, and so on. The total number of vertex inspections is $V+(V-1)+\cdots+1=\Theta(V^{2})$, establishing a $\Omega(V^{2})$ lower bound for this single scale; hence $T(V)\geq\Omega(V^{2})$.
@@
For a graph $G$ with diameter $\mathrm{diam}(G)$, the standard error of linear regression for the Minkowski dimension satisfies
$SE(\hat{m})\sim\sqrt{6}\sigma\frac{1}{\sqrt{\mathrm{diam}(G)}\log\mathrm{diam}(G)}\,.$
@@
For a uniform distribution on the interval $[0,\log\left(\lfloor\mathrm{diam}(G)/2\rfloor\right)]$, the variance is given by
$\operatorname{Var}(x)=\frac{\left(\log\left(\lfloor\mathrm{diam}(G)/2\rfloor\right)\right)^{2}}{12}.$
If there are $n=\lfloor\mathrm{diam}(G)/2\rfloor$ data points, then the total sum of squares about the mean is
$S_{xx}\ =\lfloor\mathrm{diam}(G)/2\rfloor\cdot\frac{\left(\log\left(\lfloor\mathrm{diam}(G)/2\rfloor\right)\right)^{2}}{12}.$
Let $\sigma$ be the residual standard deviation. We obtain
$\operatorname{SE}(\hat{m})=\sigma\sqrt{\frac{12}{\lfloor\mathrm{diam}(G)/2\rfloor\,\left(\log\left(\lfloor\mathrm{diam}(G)/2\rfloor\right)\right)^{2}}}$
As a result,
$\lim_{n\to\infty}\frac{SE(\hat{m})}{\big{(}\sqrt{\mathrm{diam}(G)}\log\mathrm{diam}(G)\big{)}^{-1}}=\lim_{n\to\infty}\frac{\sqrt{12}\sigma}{\sqrt{2}}\frac{1}{\sqrt{\mathrm{diam}(G)}(\log\mathrm{diam}(G)-\log 2)}=\sqrt{6}\sigma$
This completes the proof. ∎
@@
    Residuals. For each scale $l$, the log–regression errors $\varepsilon_{l}$ (on $G$) and $\varepsilon^{\mathscr{R}}_{l}$ (on $\mathscr{R}(G)$) are independent, centred, and share the same finite variance $\sigma^{2}$ and finite fourth moment.
@@
    Design matrix growth. With $x_{i}=\log(1/r_{i})$ and $n=\lfloor\operatorname{diam}(G)/2\rfloor$, we have $\frac{1}{n}\sum_{i=1}^{n}x_{i}=0$ and $\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}\to\mathbb{E}[X^{2}]$ as $n\sim\operatorname{diam}(G)/2\to\infty$.
@@
    Cross-graph independence. The two residual sequences $\{\varepsilon_{l}\}_{l}$ and $\{\varepsilon^{\mathscr{R}}_{l}\}_{l}$ are mutually independent (or weakly dependent in a way that preserves the OLS CLT).
@@
Let $\mu_{G}$ be the probability measure induced by the random variable $\Delta(G)=\dim(G)-\dim\bigl{(}\mathscr{R}(G)\bigr{)}$ on a graph $G$ of diameter $D=\operatorname{diam}(G)$. Under Hypothesis A.6 and with the notation of Lemmas 3.73.8, we have
[…]
as $D\to\infty$. In particular, $\tau^{2}(D)\to 0$, so the limiting distribution degenerates to the Dirac measure $\delta_{0}$; i.e. $\Delta(G)\to 0$ in probability.
@@
[…]
@@
Subtracting the two Gaussian limits and dividing by $\sqrt{D}$ gives
[…]
Multiplying and dividing by the common factor $S_{xx}^{-1}(G)\sim 6/[D(\log D)^{2}]$ from Lemma 3.7, and noting that $S_{xx}(G)\sim S_{xx}\bigl{(}\mathscr{R}(G)\bigr{)}$, we obtain the variance term in the statement,
$\tau^{2}(D)=\frac{\sigma^{2}}{S_{xx}(G)}+\frac{\sigma^{2}}{S_{xx}\bigl{(}\mathscr{R}(G)\bigr{)}}=6\sigma^{2}\,[\,D(\log D)^{2}]^{-1}.$
@@
Assumption (A4) together with Theorem 3.3 implies $m_{\mathscr{R}}-m_{G}\to 0$ as $D\to\infty$. In (A) this term is therefore negligible relative to the $D^{-1/2}$–scaled Gaussian component. Slutsky’s theorem hence yields the weak convergence
[…]
@@
[…]
@@
[…]
so the dimension discrepancy vanishes in probability. Equivalently, because the Gaussian law $\mathcal{N}\bigl{(}0,\tau^{2}(D)\bigr{)}$ has variance $\tau^{2}(D)\to 0$, its distribution $\mu_{\Delta(G)}$ weakly converges to the Dirac measure $\delta_{0}$:
@@
The main theorem shows that, with uniform log–scale sampling and i.i.d. Gaussian residuals, the gap $\Delta(G)=\dim(G)-\dim\bigl{(}\mathscr{R}(G)\bigr{)}$ is asymptotically normal, $\Delta(G)\sim\mathcal{N}\bigl{(}0,\tau^{2}(D)\bigr{)}$ with
$\tau^{2}(\mathrm{diam}(G))=\frac{6\hat{\sigma}^{2}}{\mathrm{diam}(G)\,[\log\mathrm{diam}(G)]^{2}}\;\longrightarrow 0$
so the two dimensions become statistically indistinguishable as $D\to\infty$.
@@ B.1 Preliminary Experiments
[…]
@@
REDDIT-MULTI-5K $3$$\pm$ 1.58 38.19 $\pm$ 1.71 +4.86 $2.25\times 10^{-4}$
Appendix C Experimental Methodology and Results of FractalGCL on Urban Districts
@@ C.2 Hyperparameter Configuration
C.3 Traffic Accident Classification Tasks
@@
Binary classification: compute Shannon entropy of severity‐level proportions $\{p_{i}\}_{i=1}^{4}$, then label = 1 if entropy $>$ median, else 0. Measures embedding of severity diversity.
@@
$\text{label}=\begin{cases}2&\text{if volume $>$ median \emph{and} ($sev3+sev4$)/total $>$ median},\\ 0&\text{if volume $\leq$ median \emph{and} ($sev3+sev4$)/total $\leq$ median},\\ 1&\text{otherwise}.\end{cases}$
@@
Task City DGI InfoGraph GCL JOAO SimGRACE FractalGCL
total_accidents_high Chicago 54.91$\pm$10.75 56.54$\pm$11.85 63.12$\pm$13.36 55.86$\pm$12.21 62.75$\pm$13.50 64.60$\pm$13.32
SF 76.45$\pm$14.47 78.74$\pm$13.60 80.06$\pm$13.32 79.75$\pm$13.61 80.40$\pm$13.47 80.89$\pm$12.92
NY 51.51$\pm$7.62 51.85$\pm$8.82 55.83$\pm$12.38 51.10$\pm$11.01 52.27$\pm$12.06 68.39$\pm$13.84
accident_volume_level Chicago 43.09$\pm$11.96 42.77$\pm$12.15 46.90$\pm$13.71 43.15$\pm$12.55 46.26$\pm$13.34 48.83$\pm$13.68
SF 55.85$\pm$11.48 58.07$\pm$10.67 58.31$\pm$10.86 58.40$\pm$10.61 58.43$\pm$11.10 58.10$\pm$11.37
NY 37.14$\pm$10.59 35.72$\pm$10.21 39.22$\pm$13.34 35.86$\pm$12.14 36.01$\pm$13.09 50.17$\pm$14.79
severity_entropy Chicago 50.99$\pm$7.10 52.64$\pm$9.11 58.87$\pm$12.33 52.38$\pm$9.60 57.07$\pm$11.97 65.60$\pm$13.22
SF 49.04$\pm$4.75 49.03$\pm$6.39 48.86$\pm$9.93 48.98$\pm$11.00 48.96$\pm$9.93 49.15$\pm$10.60
NY 52.23$\pm$8.21 51.86$\pm$8.61 54.45$\pm$11.82 51.63$\pm$11.03 52.40$\pm$12.00 61.27$\pm$13.49
has_sev3 Chicago 80.09$\pm$18.52 79.25$\pm$17.82 80.82$\pm$14.07 79.65$\pm$16.10 80.79$\pm$14.96 80.48$\pm$14.16
SF 98.04$\pm$0.31 97.50$\pm$6.45 97.81$\pm$3.73 97.11$\pm$4.27 97.82$\pm$3.62 97.66$\pm$3.84
NY 55.68$\pm$14.73 55.96$\pm$14.92 57.37$\pm$14.83 54.69$\pm$15.12 55.75$\pm$15.25 64.82$\pm$13.53
has_sev4 Chicago 53.90$\pm$10.67 54.00$\pm$10.17 60.37$\pm$13.48 54.02$\pm$11.12 59.76$\pm$13.76 62.74$\pm$13.66
SF 54.73$\pm$13.45 53.91$\pm$14.23 55.50$\pm$13.15 54.95$\pm$13.75 55.40$\pm$13.71 56.44$\pm$14.51
NY 52.44$\pm$11.64 53.33$\pm$12.15 56.22$\pm$13.55 53.34$\pm$13.71 54.89$\pm$14.07 59.88$\pm$13.23
risk_level Chicago 34.19$\pm$7.29 34.02$\pm$7.68 40.49$\pm$12.05 33.90$\pm$9.30 37.74$\pm$11.12 43.95$\pm$12.63
SF 39.47$\pm$11.15 40.98$\pm$10.84 41.26$\pm$11.59 41.76$\pm$12.48 41.41$\pm$11.76 42.34$\pm$12.20
NY 41.46$\pm$11.21 42.06$\pm$11.21 47.30$\pm$13.63 41.40$\pm$12.38 42.80$\pm$13.05 57.24$\pm$14.23
Average ( % ) 54.51 54.90 57.93 54.89 56.72 61.81
C.4 Urban Feature Classification Tasks
We define six downstream classification tasks based solely on the seven non‐accident features (population density and six POI densities):
dominant_poi (6‐class)
Identify the single most abundant POI category (office, sustenance, transportation, retail, leisure, residence) in each catchment.
poi_entropy_level (3‐class)
Compute the Shannon entropy of the six POI density proportions, then split into Low/Mid/High mixture levels via tertile cut.
pop_density_level (4‐class)
Partition population density into four quartile bins (Suburban, Inner Suburban, Urban, Core) labeled 0–3.
func_density_combo (12‐class)
Combine the 6 “dominant POI” classes with a binary high/low population density flag to form 12 composite categories.
nightlife_hotspot (binary)
Label catchments as nightlife hotspots (1) if the leisure‐to‐residence POI ratio exceeds its median, else non‐hotspot (0).
commercial_residential_mix (3‐class)
Compute the retail/(retail+residence) ratio and split into Residential‐Dominant (0), Mixed (1), and Commercial‐Dominant (2) via tertile cut.
Pleae find full results in Table 8.
Table 8: Urban POI–related tasks (%, mean±$\pm$±std). Bold indicates the best score in each city.
Task City DGI InfoGraph GCL JOAO SimGRACE FractalGCL
dominant_poi Chicago
SF 70.74$\pm$16.14 73.13$\pm$16.27 74.23$\pm$12.35 70.88$\pm$15.31 73.97$\pm$12.97 80.75$\pm$12.48
NY 45.31$\pm$15.49 47.09$\pm$14.85 49.97$\pm$14.56 47.08$\pm$15.78 48.27$\pm$14.97 49.42$\pm$14.90
poi_entropy_level Chicago 36.33$\pm$9.39 38.88$\pm$10.51 38.08$\pm$12.14 35.72$\pm$10.20 38.58$\pm$11.71 39.79$\pm$12.32
SF 40.21$\pm$11.09 39.30$\pm$11.13 40.00$\pm$11.77 40.01$\pm$12.86 40.48$\pm$12.21 45.48$\pm$13.40
NY 37.49$\pm$10.81 36.71$\pm$10.28 37.60$\pm$12.54 35.83$\pm$11.82 36.99$\pm$12.93 40.59$\pm$14.02
pop_density_level Chicago 32.91$\pm$11.00 36.42$\pm$12.07 36.97$\pm$12.75 33.11$\pm$11.41 37.44$\pm$12.75 39.55$\pm$13.13
SF 47.16$\pm$12.78 49.63$\pm$12.79 50.78$\pm$12.83 46.98$\pm$12.57 51.06$\pm$12.73 51.41$\pm$12.80
NY 40.09$\pm$12.61 41.35$\pm$12.77 42.17$\pm$13.06 39.77$\pm$12.72 40.75$\pm$13.42 42.76$\pm$13.96
func_density_combo Chicago
SF 50.86$\pm$13.54 53.19$\pm$12.90 55.41$\pm$12.07 51.96$\pm$13.86 55.52$\pm$11.98 58.63$\pm$12.46
NY 32.41$\pm$13.14 33.18$\pm$12.95 34.84$\pm$12.43 33.11$\pm$13.13 34.46$\pm$12.58 34.73$\pm$12.31
nightlife_hotspot Chicago 52.97$\pm$8.83 55.26$\pm$10.68 54.33$\pm$11.60 53.30$\pm$10.04 55.20$\pm$11.67 56.06$\pm$11.90
SF 54.81$\pm$10.20 55.72$\pm$11.26 57.00$\pm$12.61 53.32$\pm$12.23 56.07$\pm$11.95 61.48$\pm$13.45
NY 54.89$\pm$9.96 55.15$\pm$10.86 55.27$\pm$12.74 54.03$\pm$12.47 55.24$\pm$13.19 54.16$\pm$13.48
commercial_residential_mix Chicago 37.48$\pm$9.88 37.66$\pm$9.93 36.98$\pm$11.48 36.93$\pm$10.52 37.05$\pm$11.26 36.59$\pm$11.71
SF 50.14$\pm$13.74 50.35$\pm$14.31 52.74$\pm$14.27 49.61$\pm$13.81 52.94$\pm$14.32 53.01$\pm$14.15
NY 37.30$\pm$10.69 37.17$\pm$10.95 36.96$\pm$12.65 35.93$\pm$11.98 36.53$\pm$12.90 35.92$\pm$13.17
Average 45.07 46.26 47.08 44.85 46.91 48.77
C.5 Conclusion

### v3

@@ Appendix A Proofs
For $i\in\{1,2\}$, write
$s(H_{i}):=\mathrm{sim}\bigl(\mathbf{z}_{G},\mathbf{z}_{H_{i}}\bigr),\qquad\Delta(H_{i}):=\dim_{B}(H_{i})-\dim_{B}\!\bigl(\mathscr{R}(H_{i})\bigr),\qquad w(H_{i}):=\exp\!\bigl(\alpha\Delta(H_{i})\bigr).$
The single-sample fractal-weighted InfoNCE loss for candidate $H_{i}$ is
$\ell_{\mathrm{fractal}}(G,H_{i})=-\frac{s(H_{i})}{\tau}-\log w(H_{i})+\log Z,\qquad i=1,2,$
where $Z$ is the common partition term. Taking the difference gives
$\ell_{\mathrm{fractal}}(G,H_{2})-\ell_{\mathrm{fractal}}(G,H_{1})=-\frac{s(H_{2})-s(H_{1})}{\tau}-\log\frac{w(H_{2})}{w(H_{1})}.$
Since $\Delta(H_{2})>\Delta(H_{1})$ and $w(H)=\exp(\alpha\Delta(H))$,
$\log\frac{w(H_{2})}{w(H_{1})}=\alpha\bigl(\Delta(H_{2})-\Delta(H_{1})\bigr)>0.$
By the assumption $s(H_{1})-s(H_{2})\leq\tau\alpha\bigl(\Delta(H_{2})-\Delta(H_{1})\bigr)$, we have
$\frac{s(H_{2})-s(H_{1})}{\tau}\;\leq\;\alpha\bigl(\Delta(H_{2})-\Delta(H_{1})\bigr).$
Therefore
$\ell_{\mathrm{fractal}}(G,H_{2})-\ell_{\mathrm{fractal}}(G,H_{1})\;\leq\;0,$
and the inequality is strict whenever $s(H_{1})-s(H_{2})<\tau\alpha\bigl(\Delta(H_{2})-\Delta(H_{1})\bigr)$. ∎
@@
Lower bound $\boldsymbol{\Omega(V^{2})}$. Consider a path of $V$ vertices. At the smallest scale (covering radius $1$) a single box covers at most two vertices, so roughly $V/2$ boxes must be chosen. Each choice is made by scanning all currently uncovered vertices to find the one whose radius-$1$ neighbourhood is largest: first $V$ scans, then V1V-1, and so on. The total number of vertex inspections is $V+(V-1)+\cdots+1=\Theta(V^{2})$, establishing a $\Omega(V^{2})$ lower bound for this single scale; hence $T(V)\geq\Omega(V^{2})$.
@@
Let $G$ be a graph with diameter $\mathrm{diam}(G)$ and let $L_{\max}:=\lfloor\mathrm{diam}(G)/2\rfloor$. Consider the ordinary least squares fit of $y_{\ell}=\log N_{\ell}(G)$ on $x_{\ell}=\log\ell$ over admissible scales $\ell\in\{1,2,\dots,L_{\max}\}$, and let $\sigma^{2}$ be the log–residual variance. If the design points $\{x_{\ell}\}$ are (approximately) uniformly spaced over $0,\log L_{\max}]$ (i.e. log-uniform scale selection), then
$\operatorname{SE}(\hat{m})\;\sim\;\frac{2\sqrt{6}\sigma}{\sqrt{\mathrm{diam}(G)}\log\mathrm{diam}(G)}.$
@@
Write $n:=L_{\max}=\lfloor\mathrm{diam}(G)/2\rfloor$ for the number of scale points and $S_{xx}:=\sum_{\ell=1}^{L_{\max}}(x_{\ell}-\bar{x})^{2}$. Under the stated (approximate) uniformity of $x$ on $0,\log L_{\max}]$,
$\operatorname{Var}(x)\;=\;\frac{\bigl(\log L_{\max}\bigr)^{2}}{12},\qquad S_{xx}\;\approx\;n\operatorname{Var}(x)\;=\;\frac{L_{\max}\bigl(\log L_{\max}\bigr)^{2}}{12}.$
Hence the OLS slope standard error satisfies
$\operatorname{SE}(\hat{m})\;=\;\frac{\sigma}{\sqrt{S_{xx}}}\;\sim\;\sigma\sqrt{\frac{12}{L_{\max}\bigl(\log L_{\max}\bigr)^{2}}}\;=\;\frac{\sqrt{12}\sigma}{\sqrt{\lfloor\mathrm{diam}(G)/2\rfloor}\log\!\bigl(\lfloor\mathrm{diam}(G)/2\rfloor\bigr)}.$
Since $L_{\max}\sim\mathrm{diam}(G)/2$ and $\log L_{\max}=\log\mathrm{diam}(G)-\log 2$,
$\operatorname{SE}(\hat{m})\cdot\sqrt{\mathrm{diam}(G)}\log\mathrm{diam}(G)\;\longrightarrow\;2\sqrt{6}\sigma,$
which yields the stated asymptotic equivalence. ∎
@@
    Residuals. For each scale $\ell$, the log–regression errors $\{\varepsilon_{\ell}\}$ (on $G$) and $\{\varepsilon^{\mathscr{R}}_{\ell}\}$ (on $\mathscr{R}(G)$) are, within each graph, independent, centred, and share the same finite variance $\sigma^{2}$ and finite fourth moment.
@@
    Design matrix growth. With $x_{\ell}=\log\ell$ and $n=\lfloor\operatorname{diam}(G)/2\rfloor$, we have $\frac{1}{n}\sum_{\ell=1}^{n}x_{\ell}=0$ and $\frac{1}{n}\sum_{\ell=1}^{n}x_{\ell}^{2}\to\mathbb{E}[X^{2}]$ as $n\sim\operatorname{diam}(G)/2\to\infty$.
@@
    Cross-graph independence. The two residual sequences $\{\varepsilon_{\ell}\}_{\ell}$ and $\{\varepsilon^{\mathscr{R}}_{\ell}\}_{\ell}$ are mutually independent (or weakly dependent in a way that preserves the OLS CLT).
@@
Let $\mu_{G}$ be the probability measure induced by the random variable $\Delta(G)$ on a graph $G$. Under Hypothesis A.6 and with the notation of Lemmas 3.7 and 3.8, we have
$\mu_{G}\;\xrightarrow{\text{w}}\mathcal{N}\bigl(0,\kappa^{2}(\mathrm{diam}(G))\bigr),\qquad\kappa^{2}(\mathrm{diam}(G))=6\sigma^{2}[\mathrm{diam}(G)(\log\mathrm{diam}(G))^{2}]^{-1},$
as $\mathrm{diam}(G)\to\infty$. In particular, $\kappa^{2}(\mathrm{diam}(G))\to 0$, so the limiting distribution degenerates to the Dirac measure $\delta_{0}$; i.e. $\Delta(G)\to 0$ in probability.
@@
$\sqrt{\mathrm{diam}(G)}(\hat{m}_{G}-m_{G})\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),\qquad\sqrt{\mathrm{diam}(G)}(\hat{m}_{\mathscr{R}}-m_{\mathscr{R}})\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),$
@@
Subtracting the two Gaussian limits and dividing by $\sqrt{\mathrm{diam}(G)}$ gives
$\sqrt{\mathrm{diam}(G)}\bigl[(\hat{m}_{G}-m_{G})-(\hat{m}_{\mathscr{R}}-m_{\mathscr{R}})\bigr]\xrightarrow{d}\mathcal{N}\bigl(0,2\sigma^{2}\bigr).$
Multiplying and dividing by the common factor $S_{xx}^{-1}(G)\sim 6/[\mathrm{diam}(G)(\log\mathrm{diam}(G))^{2}]$ from Lemma 3.7, and noting that $S_{xx}(G)\sim S_{xx}\bigl(\mathscr{R}(G)\bigr)$, we obtain the variance term in the statement,
$\kappa^{2}(\mathrm{diam}(G))=\frac{\sigma^{2}}{S_{xx}(G)}+\frac{\sigma^{2}}{S_{xx}\bigl(\mathscr{R}(G)\bigr)}=6\sigma^{2}[\mathrm{diam}(G)(\log\mathrm{diam}(G))^{2}]^{-1}.$
@@
Assumption (A4) together with Theorem 3.3 implies $m_{\mathscr{R}}-m_{G}\to 0$ as $\mathrm{diam}(G)\to\infty$. In (A) this term is therefore negligible relative to the $\mathrm{diam}(G)^{-1/2}$–scaled Gaussian component. Slutsky’s theorem hence yields the weak convergence
$\mu_{G}\;\xrightarrow{w}\;\mathcal{N}\bigl(0,\kappa^{2}(\mathrm{diam}(G))\bigr),\qquad\kappa^{2}(D)=6\sigma^{2}[\mathrm{diam}(G)(\log\mathrm{diam}(G))^{2}]^{-1}.$
@@
Because $\kappa^{2}(\mathrm{diam}(G))\to 0$ as $\mathrm{diam}(G)\to\infty$, the normal limit collapses to the Dirac measure $\delta_{0}$, which implies $\Delta(G)\to 0$ in probability; equivalently, $\mu_{G}\xrightarrow{w}\delta_{0}$. ∎
@@
$\Delta(G)\xrightarrow{p}0\quad\Longleftrightarrow\quad\mathbf{P}\bigl(|\Delta(G)|>\varepsilon\bigr)\longrightarrow 0\quad(\forall\varepsilon>0),$
so the dimension discrepancy vanishes in probability. Equivalently, because the Gaussian law $\mathcal{N}\bigl(0,\kappa^{2}(\mathrm{diam}(G))\bigr)$ has variance $\kappa^{2}(\mathrm{diam}(G))\to 0$, its distribution $\mu_{\Delta(G)}$ weakly converges to the Dirac measure $\delta_{0}$:
@@
The main theorem shows that, with uniform log–scale sampling and i.i.d. Gaussian residuals, the gap $\Delta(G)=\dim_{\mathrm{B}}(G)-\dim_{\mathrm{B}}\bigl(\mathscr{R}(G)\bigr)$ is asymptotically normal, $\Delta(G)\sim\mathcal{N}\bigl(0,\kappa^{2}(\mathrm{diam}(G))\bigr)$ with
$\kappa^{2}(\mathrm{diam}(G))=\frac{6\hat{\sigma}^{2}}{\mathrm{diam}(G)[\log\mathrm{diam}(G)]^{2}}\;\longrightarrow 0$
so the two dimensions become statistically indistinguishable as $\mathrm{diam}(G)\to\infty$.
@@ B.1 Preliminary Experiments
    • Baseline: $\{\texttt{degree\_variance},\ \texttt{avg\_shortest\_path}\}$.
@@
REDDIT-MULTI-5K 33.33 $\pm$ 1.58 38.19 $\pm$ 1.71 +4.86 $2.25\times 10^{-4}$
Appendix C Fractality Threshold Analysis
The choice of the fractality threshold is a crucial component of our model. In Section 4.7, we have already analysed the D&D dataset under different threshold values and observed that the best performance is achieved when the threshold is set to $0.90$. To ensure rigour, we further conduct a comprehensive parameter analysis on all benchmark datasets, varying the fractality threshold in the same manner. The results are summarised in Table 7.
Table 7: Classification accuracy (10-fold CV) with different fractality thresholds.
Model NCI1 MUTAG PROTEINS D&D REDDIT-B REDDIT-M5K AVG.
GraphCL (You et al., 2020) 77.87 $\pm$ 0.41 86.80 $\pm$ 1.34 74.39 $\pm$ 0.45 78.62 $\pm$ 0.40 89.53 $\pm$ 0.84 55.99 $\pm$ 0.28 77.20 $\pm$ 0.72
FractalGCL ($R^{2}=0.70$) 78.94 $\pm$ 0.35 87.42 $\pm$ 0.92 74.65 $\pm$ 0.52 79.41 $\pm$ 0.49 89.60 $\pm$ 0.71 56.42 $\pm$ 0.44 77.74 $\pm$ 0.60
FractalGCL ($R^{2}=0.75$) 79.25 $\pm$ 0.32 88.19 $\pm$ 0.88 74.93 $\pm$ 0.47 79.88 $\pm$ 0.54 89.74 $\pm$ 0.68 56.58 $\pm$ 0.42 78.10 $\pm$ 0.58
FractalGCL ($R^{2}=0.80$) 79.62 $\pm$ 0.28 89.12 $\pm$ 0.80 75.20 $\pm$ 0.44 80.41 $\pm$ 0.52 90.00 $\pm$ 0.64 56.81 $\pm$ 0.41 78.53 $\pm$ 0.54
FractalGCL ($R^{2}=0.85$) 80.02 $\pm$ 0.22 90.20 $\pm$ 0.52 75.53 $\pm$ 0.42 81.04 $\pm$ 0.47 90.45 $\pm$ 0.61 57.02 $\pm$ 0.38 79.04 $\pm$ 0.45
FractalGCL ($R^{2}=0.90$) 80.50 $\pm$ 0.16 91.71 $\pm$ 0.23 75.85 $\pm$ 0.40 81.71 $\pm$ 0.57 90.41 $\pm$ 0.72 57.29 $\pm$ 0.59 79.58 $\pm$ 0.49
FractalGCL ($R^{2}=0.95$) 80.21 $\pm$ 0.19 91.10 $\pm$ 0.28 75.81 $\pm$ 0.36 81.22 $\pm$ 0.42 90.05 $\pm$ 0.63 57.05 $\pm$ 0.46 79.24 $\pm$ 0.41
Across most settings, $R^{2}=0.90$ emerges as the best choice. Intuitively, an excessively stringent threshold (e.g., $0.95$) excludes many graphs that are genuinely fractal, thereby reducing the proportion of samples that benefit from FractalGCL and weakening the signal. Conversely, a too–permissive threshold (e.g., 0.700.700.850.85) admits graphs whose fractality is insufficiently supported, injecting noise into the fractal loss and diluting its effectiveness. Taken together, the empirical evidence indicates that $R^{2}=0.90$ strikes the most favourable balance between coverage and reliability, and is thus our recommended default (see Table 7).
Appendix D Related Works
Fractal Geometry for Graphs. Fractal geometry interfaces with graph theory most tangibly through the study of complex networks (Watts & Strogatz, 1998; Barabási & Albert, 1999; Song et al., 2005). Analytical models of genuinely fractal graphs remain comparatively scarce. Early progress came from physicists who studied percolation and the Ising model on hierarchical lattices and Bethe trees, using real-space renormalisation to obtain non-integer critical exponents and anomalous scaling laws (Griffiths & Kaufman, 1982). Recently, the Iterated Graph Systems framework has gained attention for its rigorous yet flexible recursive construction of fractal graphs (Li & Britz, 2024; Neroli, 2024). However, applications of fractal geometry within graph representation learning are still rare.
Graph Contrastive Learning. Graph contrastive learning comprises several crucial stages, among which graph data augmentation assumes a pivotal role, yet it is rendered particularly challenging by the intricate non-Euclidean characteristics inherent in graph topologies (Ju et al., 2024). Existing graph data augmentation techniques (Velickovic et al., 2019; You et al., 2020; 2021; Qiu et al., 2020; Li et al., 2022; Wei et al., 2023; Jin et al., 2021; Ji et al., 2024) have achieved notable progress. However, they often fall short in adequately preserving the structural similarity between positive pairs, which arises from the inherent difficulty in precisely leveraging complex topological features.
Appendix E Experimental Methodology and Results of FractalGCL on Urban Districts
@@ E.2 Hyperparameter Configuration
E.3 Traffic Accident Classification Tasks
@@
Binary classification: compute Shannon entropy of severity‐level proportions $\{p_{i}\}_{i=1}^{4}$, then label = 1 if entropy is greater than median, else 0. Measures embedding of severity diversity.
@@
[…]
@@
Task City DGI InfoGraph GCL JOAO SimGRACE DRGCL GradGCL FractalGCL
[…]
E.4 Conclusion